please help me solve this !!!

(2r+1)(12r^2+5r-3)=0

Answer by ewatrrr(24785)

 
Hi, Note use of ^ (uppercase 6)
(2r+1)(12r^2 +5r-3)=0 |completing the factoring
(2r+1)(4r +3)(3r -1) = 0
(2r+1) = 0
r = -1/2
(4r +3)= 0
r = -3/4
(3r -1)= 0
r = 1/3